OpenStudy (anonymous):

For which operations is the set {–1, 1} closed? Choose all answers that are correct. A. addition B. subtraction C. multiplication D. division

OpenStudy (anonymous):

Ill give a medal

OpenStudy (anonymous):

@TuringTest

OpenStudy (perl):

Closure (under an operation) means that if you take any two things in the set, you operate on them, the result will also be in the set. So for example, if we start with the set of positive integers, N = { 1,2,3,4,5,... } , and pick two arbitrary elements, add them, the result is also in the set. Example : 2 ,3 2+3 = 5 , and 5 is in N. What about subtraction? Is the set closed {1,2,3,4... } closed under subtraction? Is 3 - 5 in the set?

OpenStudy (anonymous):

i think the answer might be subtraction but im not sure if thats right or if there are any others..

OpenStudy (perl):

it is not closed under subtraction. (-1) -1 = -2 , and -2 is not in the set

OpenStudy (perl):

how about multiplication

OpenStudy (anonymous):

that could be right

OpenStudy (perl):

1*1 = 1 , 1 is in {-1,1}. ✔ 1* (-1) = -1 , -1 is in {-1,1} . ✔ (-1)*(-1)= 1 , 1 is in {-1,1} . ✔

OpenStudy (perl):

that covers all the cases. now try division, go through the cases

OpenStudy (anonymous):

its confusing

OpenStudy (perl):

1 ÷ 1= 1 ✔ (-1) ÷ 1= - 1 ✔ (-1) ÷ (-1)= 1 ✔

OpenStudy (perl):

the result of any division is itself either 1 or -1 , a member of the set from which it originated